15. Determine the output of th ollowing Ct+ code? int num346, sum 0; do wA sum-sum + num % 10; num = num / 10; cout < sum endl; while (num> e); rmine the output produced by the following code fragment? int i 11; i 1 i 4)10 9 cout << i << . i+t; 13 cout << endl;

Please show all steps for #15 and #16 neatly.

Thank you so much in advance!

Will rate thumbs up for correct answer ! ?
Answer #1

15.

Output

6
10
13

Explanation

It is to sum all the digits in num

Each time adding last digit of num to sum

Removing last digit of the num

Doing above as long as digits are in n

16.

Output

11 8 5 2

Explanation

Starting from 11

each time decrementing i by 4 in the for line each time incrementing by 1 in the loop

Together, decrementing by 3

11 - 3 = 8

8 - 3 = 5

5 - 3 = 2